# Provably Extracting the Features from a General Superposition

### v1

@@ 6 Frequency Finding Algorithm
$f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$
@@ 6.1 Location of Nonzero Frequencies
Next, we prove two statements, Lemma 6.2 and Lemma 6.3, which characterize the $v,A$ for which the Fourier mass $I^{*}_{f^{(\ell)}}(v,A)$ can be non-negligible. First, we show that when $v$ is far (in the norm induced by $A$) from every hidden direction line $\{tv_{i}:t\in{\mathbb{R}}\}$, then the Fourier mass is small. To interpret the bound in Lemma 6.2, we will set $A$ so that $\left\lVert A\right\rVert\leq\ell^{2}$, so then the Fourier mass $I^{*}_{f^{(\ell)}}(v,A)$ decays exponentially in $D^{2}$ where $D$ is the $A$-distance from $v$ to the union of lines $\{tv_{i}:t\in{\mathbb{R}}\}_{i\in[n]}$.
@@
Let $f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying $|a_{i}|\leq 1,\left\lVert\sigma_{i}\right\rVert_{\infty}\leq 1\;\forall i$. Let $A\succeq 0$ be any positive semidefinite matrix. Define the $A$-distance from $v$ to the union of lines by
@@
We use the shorthand $I^{*}\mathrel{\mathop{:}}=I^{*}_{f^{(\ell)}}(v,A)$. Using 4.11 and Cauchy–Schwarz,
@@
$\begin{split}I^{*}\;\leq\;n\sum_{i=1}^{n}\pi^{\frac{d-1}{2}}\,\ell^{d-1}\!\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}\exp\!\Big(-\,\frac{\ell^{2}}{\ell^{2}+\|A\|}\;(v-tv_{i})^{\top}A(v-tv_{i})\Big)dt\\ \;\leq\;n\sum_{i=1}^{n}\pi^{\frac{d-1}{2}}\,\ell^{d-1}\,e^{-\frac{\ell^{2}}{\ell^{2}+\|A\|}D_{i}^{2}}\!\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}dt.\end{split}$
@@
Next, we prove a counterpart to Lemma 6.2, showing that for each direction $v_{i}$ with a non-degenerate activation $\sigma_{i}$, there is a bounded scale $\beta$ with $\beta$ also bounded away from 0 such that if $v$ is close to $\beta v_{i}$ and $A$ is not too large, then the Fourier mass $I^{*}_{f^{(\ell)}}(v,A)$ is bounded away from zero.
@@
Let $f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying $|a_{i}|\leq 1,\left\lVert\sigma_{i}\right\rVert_{\infty}\leq 1\;\forall i$. Assume the directions are $\gamma$-separated i.e. the sines of all angles between them are at least $\gamma$. Fix $i\in[n]$ and assume $\sigma_{i}$ is $L$-Lipschitz and $(R,\Delta)$-nondegenerate where $R\geq 1,\Delta<1$. Let $\ell\geq 20(R+L)/\Delta$, and define
@@
$\begin{split}I^{*}_{f^{(\ell)}}(v,A)\;\geq\;\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\Bigg[\;\frac{\Delta^{2}}{64n(1+B\sqrt{\alpha})R\,\ell^{2}}\;-\;4n\ell\;\exp\!\Big(-\,\frac{\alpha\ell^{2}\gamma^{2}a^{2}}{\ell^{2}+\alpha}\Big)\;\Bigg]\,.\end{split}$ (11)
@@
$I^{*}_{j}\;\mathrel{\mathop{:}}=\;\int_{{\mathbb{R}}^{d}}\Big|\widehat{\sigma_{j}^{(\ell)}}(v_{j}^{\top}y)\Big|^{2}\,\ell^{2d-2}\,e^{-\ell^{2}\|y-(v_{j}^{\top}y)v_{j}\|^{2}}\,e^{-(y-v)^{\top}A(y-v)}dy.$
We first lower bound the single-component contribution $I^{*}_{i}$. Set $v=\beta v_{i}$ and decompose $y=tv_{i}+z$ with $z\in v_{i}^{\perp}$; then $\langle v,v_{i}\rangle=\beta$. Recall $A=\alpha I_{d}$. Now we follow a similar calculation to Lemma 6.2 where we first integrate over $z\in v^{\perp}$ and then over $t$ to get
$I^{*}_{i}=\int_{{\mathbb{R}}}|\widehat{\sigma_{i}^{(\ell)}}(t)|^{2}\;\ell^{2d-2}\int_{v_{i}^{\perp}}e^{-(\ell^{2}+\alpha)\left\lVert z\right\rVert^{2}}e^{-\alpha(t-\beta)^{2}}dzdt=\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}e^{-\alpha(t-\beta)^{2}}dt.$
@@
$I^{*}_{i}\;\geq\;\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\frac{\Delta^{2}}{64(1+B\sqrt{\alpha})R\,\ell^{2}}\,.$
It remains to control the remainder $\sum_{j\neq i}I^{*}_{j}$ in (12). Consider a fixed index $j$. We apply the same approach of integrating over $v_{j}^{\perp}$ and then over $t$. We set $A=\alpha I_{d}$ and $v=\beta v_{i}$ and also let $w_{j}=P_{v_{j}^{\perp}}v$ and get
$\begin{split}I^{*}_{j}&=\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}\;\ell^{2d-2}\int_{v_{i}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert tv_{j}+z-v\right\rVert^{2}}dzdt\\ &=\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}\int_{v_{j}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dzdt\\ &=\left(\int_{v_{j}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}dz\right)\left(\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dt\right)\\ &=\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\min_{z\in v_{j}^{\perp}}\left(e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}\right)\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dt\\ &\leq\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\exp\!\Big(-\,\frac{\alpha\,\ell^{2}}{\ell^{2}+\alpha}\,\|w_{j}\|^{2}\Big)\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}dt\\ &\leq\frac{\pi^{\frac{d}{2}}\,\ell^{2d-1}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\exp\!\Big(-\,\frac{\alpha\,\ell^{2}}{\ell^{2}+\alpha}\,\|w_{j}\|^{2}\Big)\end{split}$
@@
$\sum_{j\neq i}I^{*}_{j}\;\leq\;(n-1)\,\frac{\pi^{\frac{d}{2}}\,\ell^{2d-1}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\exp\!\Big(-\,\frac{\alpha\ell^{2}\gamma^{2}a^{2}}{\ell^{2}+\alpha}\Big).$
@@ 17   Recurse on $\alpha_{1},\dots,\alpha_{k+1}$
$f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$
@@
We say a point $v\in{\mathbb{R}}^{d}$ is heavy if $I^{*}_{f^{(\ell)}}(v,4C_{2}I_{d})\geq 10^{3}\varepsilon(\pi\ell^{2})^{d/2}$.
@@
Assume that $f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ is a sum of $n$ features satisfying 1 and 2. Then with parameters as in (13) and assuming $d\geq 3,R>1$ and
@@ 7 Function Recovery
Once we have recovered the directions that are close to the $v_{i}$, we now show how to recover the actual functions $\sigma_{i}$ and the corresponding coefficients $a_{i}$. An important subroutine for this step is using queries to estimate the value of $\widehat{f^{(\ell)}}(y)$ at any specified point $y$.
@@ 5      Set $c_{j}\leftarrow f_{\sim}(x)e^{-\mathrm{i}y^{\top}x}$
Now our next goal will be to show that for an unknown sum of features $f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ satisfying 1 and 2, if we are given a direction $u$ that is sufficiently close to $v_{i}$ for some $i\in[n]$, then we can recover the corresponding function $\sigma_{i}$. First, we begin by setting parameters We assume we are given parameters $d,n,R,L,\gamma,\varepsilon$ and $\varepsilon\leq\frac{1}{\mathrm{poly}(d,n,R,L,\frac{1}{\gamma})}$ for some sufficiently large polynomial. We then set:
@@
Let $f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying 1 and 2. Assume that we are given a unit vector $u$ such that $\left\lVert u-v_{i}\right\rVert\leq\Delta$ (with parameters set in (14)). Then for any $t$ with $\frac{1}{\ell^{0.9}}\leq|t|\leq\frac{1}{\Delta^{0.1}}$,
$|\widehat{f^{(\ell)}}(tu)-a_{i}\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)|\leq 2\Delta^{0.8}\ell^{d}\,.$
@@
$\widehat{f^{(\ell)}}(tv_{i})=\sum_{j=1}^{n}a_{j}\widehat{\sigma_{j}^{(\ell)}}(tv_{j}^{\top}v_{i})\cdot\ell^{d-1}e^{-\ell^{2}\left\lVert tv_{i}-(tv_{j}^{\top}v_{i})v_{j}\right\rVert^{2}/2}$
@@
Thus, by the way we set $\ell$, we can ensure the sum of all of the terms for $j\neq i$ is at most $\Delta^{0.8}\ell^{d}$. The term for $j=i$ is exactly $a_{i}\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)$ so putting these together, we get
$|\widehat{f^{(\ell)}}(tu)-a_{i}\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)|\leq 2\Delta^{0.8}\ell^{d}$
@@
Let $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be a sum of features $f=a_{1}\sigma_{1}(v_{1}^{\top}x_{1})+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)$ satisfying 1 and 2. Suppose we are given a unit vector $u$ such that there exists some $i\in[n]$ with $\left\lVert u-v_{i}\right\rVert\leq\Delta$. Then with parameters set as in (14), there is an algorithm that takes $\mathrm{poly}(1/\Delta)$ queries to a Fourier value oracle $V_{\tau,f^{(\ell)}}$ and runtime and outputs a function $\widetilde{\sigma}$$\mathbb{R}$$\to{\mathbb{R}}$ with
$\max_{x\in{\mathbb{R}}^{d},\left\lVert x\right\rVert\leq R}|\widetilde{\sigma}(u^{\top}x)-a_{i}\sigma_{i}(v_{i}^{\top}x)|\leq\frac{5}{\ell^{0.7}}\,.$
@@
$|a_{i}\sigma_{i}(z)-\phi_{0}(z)|\leq\Delta^{0.04}+\frac{2}{\Delta^{0.1}}\cdot 2\Delta\ell^{2}\leq\Delta^{0.03}\,.$
@@
$|\widetilde{\sigma}(z)-a_{i}\sigma_{i}(z)|\leq\Delta^{0.6}+\Delta^{0.03}+\frac{3}{\ell^{0.7}}\leq\frac{4}{\ell^{0.7}}\,.$
@@
$|a_{i}\sigma_{i}(v_{i}^{\top}x)-a_{i}\sigma_{i}(u^{\top}x)|\leq R\Delta\ell^{2}\leq\frac{1}{\ell^{0.7}}\,.$
@@
$|\widetilde{\sigma}(u^{\top}x)-a_{i}\sigma_{i}(v_{i}^{\top}x)|\leq|\widetilde{\sigma}(u^{\top}x)-a_{i}\sigma_{i}(u^{\top}x)|+|a_{i}\sigma_{i}(v_{i}^{\top}x)-a_{i}\sigma_{i}(u^{\top}x)|\leq\frac{5}{\ell^{0.7}}$

### v2

@@ 6 Frequency Finding Algorithm
$f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$
@@ 6.1 Location of Nonzero Frequencies
Next, we prove two statements, Lemma 6.2 and Lemma 6.3, which characterize the $v,A$ for which the Fourier mass $I^{\star}_{f^{(\ell)}}(v,A)$ can be non-negligible. First, we show that when $v$ is far (in the norm induced by $A$) from every hidden direction line $\{tv_{i}:t\in{\mathbb{R}}\}$, then the Fourier mass is small. To interpret the bound in Lemma 6.2, we will set $A$ so that $\left\lVert A\right\rVert\leq\ell^{2}$, so then the Fourier mass $I^{\star}_{f^{(\ell)}}(v,A)$ decays exponentially in $D^{2}$ where $D$ is the $A$-distance from $v$ to the union of lines $\{tv_{i}:t\in{\mathbb{R}}\}_{i\in[n]}$.
@@
Let $f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying $\left\lVert\sigma_{i}\right\rVert_{\infty}\leq 1\;\forall i$. Let $A\succeq 0$ be any positive semidefinite matrix. Define the $A$-distance from $v$ to the union of lines by
@@
We use the shorthand $I^{\star}\mathrel{\mathop{:}}=I^{\star}_{f^{(\ell)}}(v,A)$. Using 4.11 and Cauchy–Schwarz,
@@
$\begin{split}I^{\star}\;\leq\;n\sum_{i=1}^{n}\pi^{\frac{d-1}{2}}\,\ell^{d-1}\!\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}\exp\!\Big(-\,\frac{\ell^{2}}{\ell^{2}+\|A\|}\;(v-tv_{i})^{\top}A(v-tv_{i})\Big)dt\\ \;\leq\;n\sum_{i=1}^{n}\pi^{\frac{d-1}{2}}\,\ell^{d-1}\,e^{-\frac{\ell^{2}}{\ell^{2}+\|A\|}D_{i}^{2}}\!\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}dt.\end{split}$
@@
Next, we prove a counterpart to Lemma 6.2, showing that for each direction $v_{i}$ with a non-degenerate activation $\sigma_{i}$, there is a bounded scale $\beta$ with $\beta$ also bounded away from 0 such that if $v$ is close to $\beta v_{i}$ and $A$ is not too large, then the Fourier mass $I^{\star}_{f^{(\ell)}}(v,A)$ is bounded away from zero.
@@
Let $f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying $\left\lVert\sigma_{i}\right\rVert_{\infty}\leq 1\;\forall i$. Assume the directions are $\gamma$-separated i.e. the sines of all angles between them are at least $\gamma$. Fix $i\in[n]$ and assume $\sigma_{i}$ is $L$-Lipschitz and $(R,\Delta)$-nondegenerate where $R\geq 1,\Delta<1$. Let $\ell\geq 20(R+L)/\Delta$, and define
@@
$\begin{split}I^{\star}_{f^{(\ell)}}(v,A)\;\geq\;\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\Bigg[\;\frac{\Delta^{2}}{64n(1+B\sqrt{\alpha})R\,\ell^{2}}\;-\;4n\ell\;\exp\!\Big(-\,\frac{\alpha\ell^{2}\gamma^{2}a^{2}}{\ell^{2}+\alpha}\Big)\;\Bigg]\,.\end{split}$ (12)
@@
$I^{\star}_{j}\;\mathrel{\mathop{:}}=\;\int_{{\mathbb{R}}^{d}}\Big|\widehat{\sigma_{j}^{(\ell)}}(v_{j}^{\top}y)\Big|^{2}\,\ell^{2d-2}\,e^{-\ell^{2}\|y-(v_{j}^{\top}y)v_{j}\|^{2}}\,e^{-(y-v)^{\top}A(y-v)}dy.$
We first lower bound the single-component contribution $I^{\star}_{i}$. Set $v=\beta v_{i}$ and decompose $y=tv_{i}+z$ with $z\in v_{i}^{\perp}$; then $\langle v,v_{i}\rangle=\beta$. Recall $A=\alpha I_{d}$. Now we follow a similar calculation to Lemma 6.2 where we first integrate over $z\in v^{\perp}$ and then over $t$ to get
$I^{\star}_{i}=\int_{{\mathbb{R}}}|\widehat{\sigma_{i}^{(\ell)}}(t)|^{2}\;\ell^{2d-2}\int_{v_{i}^{\perp}}e^{-(\ell^{2}+\alpha)\left\lVert z\right\rVert^{2}}e^{-\alpha(t-\beta)^{2}}dzdt=\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\int_{{\mathbb{R}}}\Big|\widehat{\sigma_{i}^{(\ell)}}(t)\Big|^{2}e^{-\alpha(t-\beta)^{2}}dt.$
@@
$I^{\star}_{i}\;\geq\;\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\frac{\Delta^{2}}{64(1+B\sqrt{\alpha})R\,\ell^{2}}\,.$
It remains to control the remainder $\sum_{j\neq i}I^{\star}_{j}$ in (13). Consider a fixed index $j$. We apply the same approach of integrating over $v_{j}^{\perp}$ and then over $t$. We set $A=\alpha I_{d}$ and $v=\beta v_{i}$ and also let $w_{j}=P_{v_{j}^{\perp}}v$ and get
$\begin{split}I^{\star}_{j}&=\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}\;\ell^{2d-2}\int_{v_{i}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert tv_{j}+z-v\right\rVert^{2}}dzdt\\ &=\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}\int_{v_{j}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dzdt\\ &=\left(\int_{v_{j}^{\perp}}e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}dz\right)\left(\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dt\right)\\ &=\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\min_{z\in v_{j}^{\perp}}\left(e^{-\ell^{2}\left\lVert z\right\rVert^{2}-\alpha\left\lVert z-w_{j}\right\rVert^{2}}\right)\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}e^{-\alpha\left\lVert tv_{j}-v+w_{j}\right\rVert^{2}}dt\\ &\leq\frac{\pi^{\frac{d-1}{2}}\,\ell^{2d-2}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\exp\!\Big(-\,\frac{\alpha\,\ell^{2}}{\ell^{2}+\alpha}\,\|w_{j}\|^{2}\Big)\int_{{\mathbb{R}}}|\widehat{\sigma_{j}^{(\ell)}}(t)|^{2}dt\\ &\leq\frac{\pi^{\frac{d}{2}}\,\ell^{2d-1}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\exp\!\Big(-\,\frac{\alpha\,\ell^{2}}{\ell^{2}+\alpha}\,\|w_{j}\|^{2}\Big)\end{split}$
@@
$\sum_{j\neq i}I^{\star}_{j}\;\leq\;(n-1)\,\frac{\pi^{\frac{d}{2}}\,\ell^{2d-1}}{(\ell^{2}+\alpha)^{\frac{d-1}{2}}}\;\exp\!\Big(-\,\frac{\alpha\ell^{2}\gamma^{2}a^{2}}{\ell^{2}+\alpha}\Big).$
@@ 17   Recurse on $\alpha_{1},\dots,\alpha_{k+1}$
$f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$
@@
We say a point $v\in{\mathbb{R}}^{d}$ is heavy if $I^{\star}_{f^{(\ell)}}(v,4C_{2}I_{d})\geq 10^{3}\varepsilon(\pi\ell^{2})^{d/2}$.
@@
Assume that $f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$ is a sum of $n$ features satisfying 1 and 2. Then with parameters as in (14) and assuming $d\geq 3,R>1$ and
@@ 7 Function Recovery
Once we have recovered the directions that are close to the $v_{i}$, we now show how to recover the actual functions $\sigma_{i}$. An important subroutine for this step is using queries to estimate the value of $\widehat{f^{(\ell)}}(y)$ at any specified point $y$.
@@ 5      Set $c_{j}\leftarrow f_{\sim}(x)e^{-\mathrm{i}y^{\top}x}$
Now our next goal will be to show that for an unknown sum of features $f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$ satisfying 1 and 2, if we are given a direction $u$ that is sufficiently close to $v_{i}$ for some $i\in[n]$, then we can recover the corresponding function $\sigma_{i}$. First, we begin by setting parameters We assume we are given parameters $d,n,R,L,\gamma,\varepsilon$ and $\varepsilon\leq\frac{1}{\mathrm{poly}(d,n,R,L,\frac{1}{\gamma})}$ for some sufficiently large polynomial. We then set:
@@
Let $f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)$ be a sum of features satisfying 1 and 2. Assume that we are given a unit vector $u$ such that $\left\lVert u-v_{i}\right\rVert\leq\Delta$ (with parameters set in (15)). Then for any $t$ with $\frac{1}{\ell^{0.9}}\leq|t|\leq\frac{1}{\Delta^{0.1}}$,
$|\widehat{f^{(\ell)}}(tu)-\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)|\leq 2\Delta^{0.8}\ell^{d}\,.$
@@
$\widehat{f^{(\ell)}}(tv_{i})=\sum_{j=1}^{n}\widehat{\sigma_{j}^{(\ell)}}(tv_{j}^{\top}v_{i})\cdot\ell^{d-1}e^{-\ell^{2}\left\lVert tv_{i}-(tv_{j}^{\top}v_{i})v_{j}\right\rVert^{2}/2}$
@@
Thus, by the way we set $\ell$, we can ensure the sum of all of the terms for $j\neq i$ is at most $\Delta^{0.8}\ell^{d}$. The term for $j=i$ is exactly $\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)$ so putting these together, we get
$|\widehat{f^{(\ell)}}(tu)-\ell^{d-1}\widehat{\sigma_{i}^{(\ell)}}(t)|\leq 2\Delta^{0.8}\ell^{d}$
@@
Let $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be a sum of features $f=\sigma_{1}(v_{1}^{\top}x_{1})+\dots+\sigma_{n}(v_{n}^{\top}x)$ satisfying 1 and 2. Suppose we are given a unit vector $u$ such that there exists some $i\in[n]$ with $\left\lVert u-v_{i}\right\rVert\leq\Delta$. Then with parameters set as in (15), there is an algorithm that takes $\mathrm{poly}(1/\Delta)$ queries to a Fourier value oracle $V_{\tau,f^{(\ell)}}$ and runtime and outputs a function $\widetilde{\sigma}$$\mathbb{R}$$\to{\mathbb{R}}$ with
$\max_{x\in{\mathbb{R}}^{d},\left\lVert x\right\rVert\leq R}|\widetilde{\sigma}(u^{\top}x)-\sigma_{i}(v_{i}^{\top}x)|\leq\frac{5}{\ell^{0.7}}\,.$
@@
$|\sigma_{i}(z)-\phi_{0}(z)|\leq\Delta^{0.04}+\frac{2}{\Delta^{0.1}}\cdot 2\Delta\ell^{2}\leq\Delta^{0.03}\,.$
@@
$|\widetilde{\sigma}(z)-\sigma_{i}(z)|\leq\Delta^{0.6}+\Delta^{0.03}+\frac{3}{\ell^{0.7}}\leq\frac{4}{\ell^{0.7}}\,.$
@@
$|\sigma_{i}(v_{i}^{\top}x)-\sigma_{i}(u^{\top}x)|\leq R\Delta\ell^{2}\leq\frac{1}{\ell^{0.7}}\,.$
@@
$|\widetilde{\sigma}(u^{\top}x)-\sigma_{i}(v_{i}^{\top}x)|\leq|\widetilde{\sigma}(u^{\top}x)-\sigma_{i}(u^{\top}x)|+|\sigma_{i}(v_{i}^{\top}x)-\sigma_{i}(u^{\top}x)|\leq\frac{5}{\ell^{0.7}}$